OpenStudy (anonymous):

Suppose that the height of an object shot straight up is given by h=512t-16t^2 (Here h is in seconds,) Find the max height and the time at which the object hits the ground

OpenStudy (bahrom7893):

h' = 512-32t t = 512/32 = 16

OpenStudy (bahrom7893):

max height is: 4096

OpenStudy (anonymous):

is it being shot from the ground or from somewhere else like a tower?

OpenStudy (anonymous):

I assume the group

OpenStudy (bahrom7893):

time when object hits the ground is easy..h=0 512t-16t^2 = 0 t1 = 0 512-16t=0 t = 512/16 = 32

OpenStudy (anonymous):

ground

OpenStudy (bahrom7893):

it's not 0, because at 0 it was on the ground, about to get shot into the air

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Vertex = highest point: x=-b/2a second part: Use the quadratic formula. You'll get two solutions for x. The bigger solution for x is the correct one.

OpenStudy (anonymous):

h=512t-16t^2, so h = -16t^2 +512t + 0 a = -16 b = 512 t=-b/2a Once you find your t-value for the vertex, plug it back in to your quadratic equation to find h. h = -16t^2 +512t + 0

OpenStudy (anonymous):

that helped a lot. Thanks a ton